The school park has a rectangular flower bed measuring 15 meters by 20 meters. The school administrators plan to double the area by providing a fence of uniform distance around the flower bed. Find the area and perimeter of the fence to the nearest hundredths.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the area of the fence is currently 15 * 20 = 300 square feet.
to double the area, it needs to become 600 square feet.
each side would have to be equal to sqrt(600) = 24.49489743
round to two decimal places (nearest hundredth) to get 24.50 feet on each side.
the perimeter of the fence would be equal to 4 * 24.49489743 = 97.97958971.
round to two decimal places to get 97.98.

the formulas used are:
area = length * width for the rectangle.
area = side squared for the square.
perimeter = 2 * (length + width) for the rectangle).
perimeter = 4 * side for the square.

all rounding is done at the end of each calculation.


Answer by ikleyn(52775)

.


            Tutor Theo misread the problem, so his solution is conceptually incorrect.

            I came to guide you in right direction.


The area of the garden is 15*20 = 30 square meters;

so the area inside the fence is 2*300 = 600 square meters.

Thus this part is solved: it is easy part.


Now let start working on the second question.



Lex x be the uniform distance around the flower bed.


Then the dimensions of the fence are  (15+2x) meters width and (20+2x) meters length.


The area inside the fence is then  (15+2x)*(20+2x).


From here you have this equation to determine x


    (15+2x)*(20+2x) = 2*15*20 = 600.


Reduce this equation to the standard quadratic equation form and solve using the quadratic formula.


    4x^2 + 70x - 300 = 0.


It has no solution/solutions in integer or rational numbers, so other solution methods (guessing or factoring)
do not work.


When you find the value of x, you determine the perimeters as

    (15+2x) + (20+2x) + (15+2x) + 20+2x).


-------------

To see many other similar solved problems (your potential TEMPLATES), look into the lessons
    - Problems on the area and the dimensions of a rectangle surrounded by a strip
    - Cynthia Besch wants to buy a rug for a room
    - Problems on a circular pool and a walkway around it
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Dimensions and the area of rectangles and circles and their elements".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.